Keisha takes out a loan for her college tuition from a bank that charges simple interest at an annual rate of 8.95%. Her loan is for 3500 for 11 months. assume each month is 1/12 of a year..
Find the interest that will be owed after 11 months
Assuming Keisha doesn’t make any payments,find the amount owed after11 months...
The part that has me baffled is assume each month is 1/12 of a year ....

Answer by stanbon(75887)

Keisha takes out a loan for her college tuition from a bank that charges simple interest at an annual rate of 8.95%. Her loan is for 3500 for 11 months. assume each month is 1/12 of a year..
Find the interest that will be owed after 11 months
Assuming Keisha doesn’t make any payments,find the amount owed after11 months...
----
I = P*r*t
-----
I = 3500*0.0895*(11/12) = $287.15
-----------------
Cheers,
Stan H.
---------------